Journalist Report – March 9th

Sol 25 – Let’s (pretend to) save the station!

“Behind the door, the stove was making pancakes which filled the house with a rich baked odor and the scent of maple syrup.”

– Chapter 25 of The Martian Chronicles by Ray Bradbury

Only two more Sols! Trying to make the most of them is the crew’s ambition. Each crewmember is working hard to wrap up their experiments, gathering and organizing the last bits of data. Despite our starting to feel quite tired, the inexorable approach of the end of the mission has not affected our enthusiasm. Nevertheless, a few signs that we’re nearing the end are showing through: we talk of what comes next, of our next adventures back on Earth, of our families and friends that we’re anxious to see again.

This morning, we prepared Alice, Adrien, and Jérémy for their last EVA in the Martian desert. Although we have gotten “used” to performing extravehicular activities, and are now more efficient than ever in preparing our crewmates to exit the station, Quentin, Alexandre, and I reminded each other of the importance of taking these EVAs extremely seriously. As we saw the two rovers disappear behind a hill, and as I gradually lost radio contact due to the growing distance, we realized once again how isolated and vulnerable we could feel, especially with half the crew missing from the Hab, and even if only two Sols remain. The Martian environment outside the Hab remains hostile, no matter how efficient and accustomed we have become.

After our explorers returned, we had a relaxing lunch, listening to and talking about the music Jérémy was playing on the speaker. What we had forgotten was that it would not be playing music very much longer… Just as Corentin and Quentin were wrapping up an EchoFinder session, and Adrien was returning from the GreenHab, the depressurization alarm resounded. Immediately, we gathered in the Lower Deck to begin simulating an emergency procedure, designed to identify how AI4U could help future astronauts follow a complex procedure. Two crewmembers toured the station to check “sensors” to find the source of the leak. The emergency EVA crew could then proceed to “repair” the station from the outside. Thanks to them, the Science Dome has been saved!

After the 30-minute operation, everyone returned to their tasks. Pancakes and bread were baked by Alice and Quentin, while I sorted and edited the many beautiful shots taken during this morning’s EVA. Quite foolishly, I was taking my time to write this report, more than usual, when Jérémy reminded all of us that tonight was our last evening on Mars… officially. I felt a rush of adrenaline: let us make the most of it!

EVA Report – March 9th

Crew 275 EVA Report 09Mar2023

EVA # 20

Author: Alice Chapiron, EVA Leader and Crew Scientist

Purpose of EVA: The purpose of the EVA was to explore the Special Region

Start time: 9:06

End time: 12:06


09:11: Airlock closed.

09:12: Rovers Opportunity and Spirit unplugged. They then left from the Hab at 09:14.

09:18: Rovers parked near the atmospheric instruments.

09:38: Data retrieved, and batteries changed for the Purple Air, LOAC, Mega-Ares, and Field Mill. The EVA crew is about to go back to the rovers to go to their target location.

10:00: Rovers are parked on the road to the Special Region to save the battery of the rovers

10:45: The EVA Crew arrived at the Special Region. They spotted a pick-up truck in the area so they decided to turn back.

11:30: Back to the rovers.

11:58: Arrival at the Hab.

12:49: End of the EVA.

Destination: Special Region

Coordinates: N4256000 E518500

Participants: EVA Leader: Alice Chapiron (Crew Scientist), Adrien Tison (Crew Botanist), and Jérémy Rabineau (Crew Commander)

Road(s) and routes per MDRS Map: Cow Dung Road until the rovers where at 65% of battery, then walk to the Gateway to Lith and the Special Region.

Mode of travel: Driving using Spirit and Opportunity and walking

Crew 275 EVA Report 09-03-2023

EVA # 21

Author: Alexandre VINAS (Crew Astronomer and HabCom)

Purpose of EVA: Emergency drill, simulation of a depressurization in the science dome corridor that had to be fixed, and test of AI4U, an artificial intelligence designed to help the astronauts during emergencies

Start time: 15:17

End time: 15:35


15:03 Emergency started

15:15 EVA crew is equipped for the EVA

15:17 Airlock closed

15:22 EVA crew out of the MDRS

15:23 EVA crew arrived at the place of the depressurization, located in the corridor of the Science Dome, on the right side when facing the Science Dome

15:27 “Hole” repaired (EVA crew took tools and simulated operating a repair)

15:35 EVA Crew back in the Hab

The hole was repaired in less than 30 minutes, that was our purpose.

Destination: In the vicinity of the MDRS, next to the Science Dome.

Coordinates (use UTM WGS 84): N4251000 E518200

Participants: EVA Leader: Corentin Senaux (HSO), Adrien Tison (Crew Botanist)

HAB COM: Alexandre VINAS (Crew Astronomer)

Road(s) and routes per MDRS Map: /

Mode of travel: Walking

GreenHab Report – March 9th

GreenHab Officer: Adrien Tison

Environmental control: Heater, (open door)

Average Temperature: 86.5°F

Average humidity: 18%

Hours of supplemental light: 0 hour

Daily water usage for crops: 15.50 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 46.87 gallons

Time(s) of watering for crops: 9:30 & 14:30

Changes to crops: N/A

Narrative: This morning it was our HSO who went to water plants as I was on an EVA.

It was already pretty hot in there; he then warned me about it.

I went to check on everything in the afternoon and I confirm that it was really hot. I saw that some plants, particularly tomatoes needed water (considering their dirt was quite dry).

I watered well plants and crops, this explains why there is more water used than usual.

I opened the door briefly for the temperature to drop down from 97°F to 86°F.

Apart from that, plants are looking good, the cucumber plant still amazes me.

And finally, I waited for the end of this narrative to talk about the fish because they are doing well.

These are good news, and a glimpse of hope for the end of this mission!

Harvest: N/A

Support/supplies needed: Even though the fish are doing well in their new habitat, I keep my proposition for Donald to retrieve them because it is not ideal to keep them in a small bucket.

Operations Report – March 9th

Crew 275 Operations Report 09Mar2023
SOL: 25

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• Robotic observatory

Notes on non-nominal systems: None.

Spirit rover used: Yes
• Hours: 210.1
• Beginning charge: 100%
• Ending charge: 53%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 114.0
• Beginning charge: 100%
• Ending charge: 60%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.8
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. An emergency depressurization protocol was simulated, and an emergency EVA was conducted, with Mission Support’s approval.
• WATER USE: 11.84
• Water: 330.11 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Starlink (MDRSCampus-Guest) is used during COMMS and during the day for emails and to handle the medical emergency. MDRScampus is used for experimental purpose.

Summary of suits and radios: Radio n°5 seems to be non-nominal. During today’s EVA, the crewmember with this radio had difficulties to be heard by the other crewmembers. The wire had to be unplugged, then plugged again and pushed very firmly, repeatedly, and much more than usual.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 15.50 gallons
• Heater: On
• Supplemental light: No
• Harvest: No

Summary of ScienceDome operations: The Commander conducted experimental protocols.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: The robotic observatory is non-nominal.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support:
• The lamp located on the right when entering the ScienceDome is missing a light bulb.

Sol Summary Report – March 9th

Crew 275 Sol Summary Report 09Mar2023

Sol: 25

Summary Title: Let’s (pretend to) save the station!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

We were very excited with today’s EVA, as we were planning to explore the Special Region and the weather was very good (which we are not used to!). Our routine operations on the atmospheric instruments went perfectly and we headed North towards the Special Region, leaving the rovers before the battery level was too low. About 45 minutes of walking later, we finally had a visual on the Special Region, however, we saw a pick-up truck parked just in front of it. We thus decided not to take the risk of meeting anyone and started our way back. Rather than exploration, it was more a kind of hike with spacesuits, but we enjoyed it very much anyway!

Back at the Hab, we continued yesterday’s training about the emergency protocols assisted by AI4U, our artificial intelligence tool. An alarm was set to ring at a random time during the afternoon and we all went on with our respective tasks. It was 15:05 when the alarm rang. The Crew Journalist took the lead of the operations and, as everyone was already in the Hab, we directly formed three teams. The Crew Scientist and the Crew Engineer explored all the buildings of the campus and restarted the environmental instruments to check where the simulated hole was located. In the meantime, the Crew Botanist and the Health & Safety Officer put on their spacesuits to get ready to fix the hole, while the other crewmembers kept an eye on the procedure and communicated the instructions to the other teams. After identifying a fictive hole in the tunnel leading to the Science Dome, a call was made to Mission support to request an emergency EVA. After being authorized to exit the airlock, the EVA crew successfully identified and fixed the fictive hole in the tunnel. The overall process took about 30 minutes, which is exactly the target time.

No bad news with the fish today. It looks like they are doing fine, but they are still in the very small bucket where they were placed yesterday. We would like to reiterate our suggestion for Donald to retrieve them whenever he wants so that they can be placed in a larger aquarium.

Look Ahead Plan:

We would like to perform an EVA dedicated to the disassembling of our atmospheric instruments tomorrow morning (Sol 26).

Then, in the early afternoon, we plan to perform another test of the emergency protocols with AI4U, our numerical assistant based on artificial intelligence. An emergency EVA may thus be requested to go to the immediate vicinity of the Hab at any time between 14:00 and 16:30.

We would like to exit the simulation at 17:30 and start the checkout procedure as well as the food inventory. Is there a crew coming to the MDRS directly after us? Do you confirm that we should arrive at the Day’s Inn in Grand Junction in the early afternoon on Saturday (Sol 27)?

Anomalies in work: None.

Weather: Sunny, low -2°C/28°F, high 9°C/48°F.

Crew Physical Status: One crewmember recovering from surgery after a broken collarbone.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site